# Spinor bilinears and Killing-Yano forms in generalized geometry

### v1

@@ Abstract
Spinor bilinears of generalized spinors and their properties are investigated. Generalized Killing and twistor spinor equations are considered and their relations to the equations satisfied by special types of differential forms are found. Killing equation in generalized geometry is written in terms of the generalized covariant derivative and Killing-Yano forms are described in the framework of generalized geometry. Construction of generalized Killing-Yano forms and generalized closed conformal Killing-Yano forms in terms of the spinor bilinears of generalized Killing spinors are determined.
@@ I Introduction
Generalized geometry is first defined as the extension of the tangent bundle of a manifold to the direct sum of the tangent and cotangent bundles [1, 2]. In that way, besides the metric, higher-degree forms can also be incorporated into the geometrical picture. The importance of this approach is that it can be directly used in the geometrical formulation of supergravity theories in various dimensions [3, 4]. By the definition of the generalized tangent bundle, the symmetry transformations of the structure group naturally includes the closed 2-forms (B-field) [2, 3]. So, the fluxes in supergravity theories, especially in 10-dimensional IIA and IIB theories, can be naturally described as part of the geometry by approaching them from the perspective of generalized geometry. Namely, generalized geometry gives way to describe the supergravity theories in a full geometrical framework. Besides the generalized vectors and generalized metric, one can also define generalizations of the Lie bracket in generalized geometry framework which gives the Courant algebroid structure to the generalized tangent bundle [5, 6]. The geometrical structure of supergravity theories can be analyzed more extensively by exploiting the definition of generalized G-structures and generalized spinors [7, 8, 9, 10]. The approach of generalized geometry can also be extended to exceptional cases by defining the generalized tangent bundle to include higher-degree form spaces [11, 12, 13, 14].
@@
Generalized Lie derivative in generalized geometry can be described as the Dorfman bracket of generalized vectors [3]. Generalized vectors that preserve the generalized metric with respect to the generalized Lie derivative are called generalized Killing vectors which are the generalized geometry versions of Killing vector fields corresponding to isometries of the ambient manifold [15]. Construction of generalized spinors and analysis of Killing superalgebra structures in generalized geometry can be realized by using generalized Killing vectors. In that way, the structure of supergravity backgrounds can be described by using generalized Killing superalgebra structures [15]. Killing vector fields have antisymmetric generalizations called Killing-Yano (KY) forms [16, 17] and spinor bilinears constructed out of KY forms can allow extending Killing superalgebra structures to include higher-degree KY forms [18, 19, 20]. These extended Killing superalgebra structures have potential to provide the refinements of the analysis of the structure of supergravity backgrounds. However, description of KY forms and construction of spinor bilinears in terms of KY forms in the context of generalized geometry remains an open problem.
@@
In this paper, we construct generalized KY forms in generalized geometry and analyze their relations with generalized Killing spinors. Properties of generalized spinor bilinears are determined for various cases and special spinor equations in generalized geometry are investigated. Special types of spinors such as geometric Killing spinors and twistor spinors which are related to supersymmetry generators of different supersymmetric theories are extended to the framework of generalized geometry via generalized connection and the equations satisfied by them are analyzed in terms of covariant derivative on the exterior bundle. Starting from the description of generalized Killing vectors in terms of generalized Lie derivative, their description in terms of generalized connection is obtained and generalized KY forms are defined by using this description. Relations between generalized Killing spinors and generalized KY forms are investigated and it is found that generalized KY forms and generalized closed conformal KY forms can be constructed from the bilinear forms of generalized Killing spinors. It is shown that, bilinear forms of generalized Killing spinors correspond to conformal KY forms in general.
@@
The paper is organized as follows. In Section 2, we summarize the basic formulas and definitions in generalized geometry which are used throughout the paper. Section 3 deals with the choices of spinor inner products for generalized spinors and the properties of generalized spinor bilinears constructed out of them. In Section 4, we consider generalized Killing and twistor spinors and find their description in terms of differential forms of the exterior bundle on the ambient manifold. Section 5 includes the description of generalized Killing vectors in terms of generalized connection and definition of generalized KY forms. In Section 6, relation between generalized Killing spinors and generalized KY forms are investigated. Section 7 concludes the paper.
@@ II Generalized Geometry
or we can also denote it in the matrix form as $\mathcal{X}=\left(\begin{array}[]{cc}X\\ \xi\end{array}\right)$. One can define a bilinear form <,><\,,\,>< , > on $E$ which is written for generalized vectors $\mathcal{X}=X+\xi$ and $\mathcal{Y}=Y+\eta$ as follows
$<\mathcal{X},\mathcal{Y}>=\frac{1}{2}\big{(}i_{X}\eta+i_{Y}\xi\big{)}$ (1)
where $i_{X}$ denotes the interior derivative or contraction with respect to the vector field $X$. So, for a single generalized vector, we have $<\mathcal{X},\mathcal{X}>=i_{X}\xi$. This bilinear form has signature $(n,n)$ and with the orientation defined for $E$, the structure group of $E$ reduces to $SO(n,n)$ [2]. The invariance of <,><\,,\,>< , > under $SO(n,n)$ gives shear transformations for generalized vectors. For a 2-form $B\in\Lambda^{2}M$ and a 2-vector $\beta\in\Lambda_{2}M$, we have $B$- and $\beta$-transforms of $\mathcal{X}$ as follows [2]
@@
which is the generalized metric induced by $g$ and $B$. Since $\mathcal{G}^{2}=I$, ±$\pm 1$ eigenspaces of $\mathcal{G}$ which are denoted by $V_{\pm}$ give a metric splitting of $E$
@@
From the basis vectors $\{X_{a}\}\in\Gamma TM$ and basis 1-forms $\{e^{a}\}\in\Gamma T^{*}M$ for $a=1,...,n$, we can define a basis for generalized vectors as
@@
where $\epsilon=\pm 1$ for $V_{\pm}$ and $A=1,...,2n$. $\mathcal{G}$-duals of $\mathcal{X}_{A}$ correspond to basis for generalized 1-forms
@@
Similar to the Lie bracket [,][\,,\,][ , ] of vector fields on $\Gamma TM$, we can also define a bracket operation on $E$ called Courant bracket as $[\,,\,]_{C}:\Gamma E\times\Gamma E\rightarrow\Gamma E$. For two generalized vectors $\mathcal{X}=X+\xi$ and $\mathcal{Y}=Y+\eta$, Courant bracket is defined as
$[\mathcal{X},\mathcal{Y}]_{C}=[X,Y]+\mathcal{L}_{X}\eta-\mathcal{L}_{Y}\xi-\frac{1}{2}d\left(i_{X}\eta-i_{Y}\xi\right)$ (7)
@@
for $X,Y,Z\in\Gamma TM$. The cohomology class of $H$ classifies the Courant algebroids on $M$ up to isomorphism [21, 22, 23]. In the presence of non-zero $H$, one can modifiy the Courant bracket and define the twisted Courant bracket $[\,,\,]_{H}$ as follows
$[\mathcal{X},\mathcal{Y}]_{H}=[\mathcal{X},\mathcal{Y}]_{C}-i_{X}i_{Y}H.$ (10)
$[\,,\,]_{H}$ also defines a Courant algebroid structure for $dH=0$. Another bracket operation on $E$ which satisfies the Jacobi identity but not antisymmetric is the Dorfman bracket $[\,,\,]_{D}$ and it is defined as
$[\mathcal{X},\mathcal{Y}]_{D}=[X,Y]+\mathcal{L}_{X}\eta-i_{Y}d\xi.$ (11)
Courant bracket correspond to the antisymmetrization of the Dorfman bracket and it can also be written as
$[\mathcal{X},\mathcal{Y}]_{C}=[\mathcal{X},\mathcal{Y}]_{D}-d<\mathcal{X},\mathcal{Y}>.$ (12)
@@
$[\mathbb{L}_{\mathcal{X}},\mathbb{L}_{\mathcal{Y}}]=\mathbb{L}_{[\mathcal{X},\mathcal{Y}]_{D}}$ (13)
@@
We can define a generalized connection $\mathbb{D}$ on $E$ which is compatible with the bilinear form <,><\,,\,>< , > corresponding to the linear operator
@@
However, there is a freedom in choosing a connection since $\mathbb{D}^{\prime}=\mathbb{D}+\alpha$ also defines a connection where $\alpha\in\Gamma(E^{*}\otimes o(E))$ and $o(E)$ denotes the bundle of skew-symmetric endomorphisms of $E$ w.r.t. <,><\,,\,>< , > [25, 26]. The generalized torsion 3-form $T_{\mathbb{D}}\in\Gamma(\Lambda^{3}E^{*})$ of a generalized connection $\mathbb{D}$ on $E$ is defined by
@@
In the presence of a non-zero 3-form field $H$, up to fixing of a divergence operator, a unique metric-compatible torsion-free generalized connection can be defined in the following form [25]
@@ III Inner product classes and Properties of Generalized Spinor Bilinears
which is isomorphic to $\mathbb{R}^{2^{n-1}}\oplus\mathbb{R}^{2^{n-1}}$. The spinor inner product $(\,,\,):S\times S\rightarrow\mathbb{R}$ of two generalized spinors $\phi$ and $\psi$ is written as
@@
[…]
@@
$n(\text{mod }4)$ 00 -swap $\mathbb{R}\text{-swap }\xi$ and -$\mathbb{R}\text{-sym}\oplus\mathbb{R}\text{-sym }\xi\eta$ $1$$\mathbb{R}\text{-sym}\oplus\mathbb{R}\text{-sym }\xi$ and -swap $\mathbb{R}\text{-swap }\xi\eta$ $2$swap $\mathbb{R}\text{-swap }\xi$ and -$\mathbb{R}\text{-skew}\oplus\mathbb{R}\text{-skew }\xi\eta$ $3$$\mathbb{R}\text{-skew}\oplus\mathbb{R}\text{-skew }\xi$ and -swap $\mathbb{R}\text{-swap }\xi\eta$
Since the spinor space is real, we have real inner products, sym and skew denotes the symmetric and skew inner products and swap means that inner product changes the chiral elements to each other, $\xi$ and $\xi\eta$ correspond to the relevant automorphism class. So, we have for $n=0,1(\text{mod }4)$ symmetric and for $n=2,3(\text{mod }4)$ skew inner products. Since, only the $n$-form part remains in the spinor inner product, one can see that, for $n=0(\text{mod }2)$, inner products of elements
$S^{+}\times S^{-}\text{ and }S^{-}\times S^{+}\text{ are zero }$
[…]
since the inner products of opposite chirality spinors cannot include even degree part and for $n=1(\text{mod }2)$, inner products of elements
$S^{+}\times S^{+}\text{ and }S^{-}\times S^{-}\text{ are zero }$
[…]
@@
$n(\text{mod }4)$ 00 -$\mathbb{R}\text{-sym}\oplus\mathbb{R}\text{-sym }\xi\eta$ $1$swap (sym) $\mathbb{R}\text{-swap (sym) }\xi\eta$ $2$$\mathbb{R}\text{-skew}\oplus\mathbb{R}\text{-skew }\xi\eta$ $3$swap (skew) $\mathbb{R}\text{-swap (skew) }\xi\eta$
So, in general, we can write the spinor inner product for generalized spinors as
@@
which corresponds to the Mukai pairing of forms. Generally, the automorphism class of the inner product is taken as $\xi$ in the literature [2]. However, it can be seen from the previous analysis of inner product classes that the automorphism class must be taken as $\xi\eta$ to have a non-zero spinor inner product. For any two generalized spinors $\phi$, $\psi$ and a generalized form $\mathcal{\alpha}$, we can write
@@
where the action of a generalized form on a generalized spinor $\alpha.\psi$ can be defined from the action of a generalized vector to a differental form given in (14) inductively. From the symmetry properties of the inner product and the choice of the automorphism class as $\xi\eta$, we have
@@
For a spinor $\phi$ and a $p$-form $\alpha_{p}$, we can define the spinor bilinear $p$-form as $(\phi,\alpha_{p}.\phi)$ and writing the spinor bilinears for a generalized spinor $\phi$ and a generalized $p$-form $\mathcal{\alpha}_{p}$, we have
@@
where we have used $\mathcal{\alpha}_{p}^{\xi\eta}=(-1)^{p(p-1)/2}(-1)^{p}\mathcal{\alpha}_{p}$. Moreover, for $n$ even, $p$ must be even, and for $n$ odd, $p$ must be odd, since for $n$ even $S^{\pm}\times S^{\pm}$ and for $n$ odd $S^{\pm}\times S^{\mp}$ inner products are non-zero. For a definite chirality spinor $\phi$, since the action of a generalized vector on a generalized spinor changes the chirality of the spinor, $\alpha_{p}.\phi$ has the same chirality if $p$ is even and has the reverse chirality if $p$ is odd. Then, by considering those properties, $(-1)^{\frac{n(n-1)+p(p+1)}{2}}=1$ only for $n+p=0(\text{mod }4)$ and
@@
is non-zero and for other cases the bilinear form is zero. In general, for two different generalized spinors $\phi$ and $\psi$ $(\phi\neq\psi)$, we have
@@
and other cases are zero. Then, only some special degree spinor bilinears are non-zero depending on $n$ and $p$. For two generalized spinors $\phi$ and $\psi$, we can write the spinor bilinear as a sum of bilinear $p$-forms in terms of generalized $p$-form basis as

### v2

@@ Abstract
Spinor bilinears of generalized geometry spinors and their properties are investigated. Generalized Killing and twistor spinor equations are considered and their relations to the equations satisfied by special types of differential forms are found. Killing equation in generalized geometry is written in terms of the generalized covariant derivative and Killing-Yano forms are described in the framework of generalized geometry. Construction of generalized Killing-Yano forms and generalized closed conformal Killing-Yano forms in terms of the spinor bilinears of generalized Killing spinors are determined.
@@ I Introduction
Generalized geometry is first defined as the extension of the tangent bundle of a manifold to the direct sum of the tangent and cotangent bundles [1, 2]. In this way, besides the metric, higher-degree forms can also be incorporated into the geometrical picture. The importance of this approach is that it can be directly used in the geometrical formulation of supergravity theories in various dimensions [3, 4]. By the definition of the generalized tangent bundle, the symmetry transformations of the structure group naturally include the closed 2-forms (B-field) [2, 3]. Therefore, the fluxes in supergravity theories, especially in 10-dimensional IIA and IIB theories, can be naturally described as part of the geometry by approaching them from the perspective of generalized geometry. Namely, generalized geometry gives way to describe the supergravity theories in a full geometrical framework. Besides the generalized vectors and generalized metric, one can also define generalizations of the Lie bracket in generalized geometry framework which give the Courant algebroid structure to the generalized tangent bundle [5, 6]. The geometrical structure of supergravity theories can be analyzed more extensively by exploiting the definition of generalized G-structures and generalized spinors [7, 8, 9, 10]. The approach of generalized geometry can also be extended to exceptional cases by defining the generalized tangent bundle to include higher-degree form spaces [11, 12, 13, 14].
@@
Generalized Lie derivative in generalized geometry can be described as the Dorfman bracket of generalized vectors [3]. Generalized vectors that preserve the generalized metric with respect to the generalized Lie derivative are called generalized Killing vectors which are the generalized geometry versions of Killing vector fields corresponding to isometries of the ambient manifold [15]. Construction of generalized geometry spinors and analysis of Killing superalgebra structures in generalized geometry can be realized by using generalized Killing vectors. In that way, the structure of supergravity backgrounds can be described by using generalized Killing superalgebra structures [15]. Killing vector fields have antisymmetric generalizations called Killing-Yano (KY) forms [16, 17] and spinor bilinears constructed out of KY forms can allow extending Killing superalgebra structures to include higher-degree KY forms [18, 19, 20, 21]. These extended Killing superalgebra structures have potential to provide the refinements of the analysis of the structure of supergravity backgrounds. However, description of KY forms and construction of spinor bilinears in terms of KY forms in the context of generalized geometry have been unexplored so far.
@@
In this paper, we construct generalized KY forms in generalized geometry and analyze their relations with generalized geometry Killing spinors. Properties of generalized spinor bilinears are determined for various cases and special spinor equations in generalized geometry are investigated. Special types of spinors such as geometric Killing spinors and twistor spinors which are related to supersymmetry generators of different supersymmetric theories are extended to the framework of generalized geometry via generalized connection and the equations satisfied by them are analyzed in terms of covariant derivative on the exterior bundle. Starting from the description of generalized Killing vectors in terms of generalized Lie derivative, their description in terms of generalized connection is obtained and generalized KY forms are defined by using this description. Relations between generalized geometry Killing spinors and generalized KY forms are investigated and it is found that generalized KY forms and generalized closed conformal KY forms can be constructed from the bilinear forms of generalized geometry Killing spinors. It is shown that bilinear forms of generalized Killing spinors correspond to conformal KY forms in general.
@@
The paper is organized as follows. In Section 2, we summarize the basic formulas and definitions in generalized geometry which are used throughout the paper. Section 3 deals with the choices of spinor inner products for generalized spinors and the properties of generalized spinor bilinears constructed out of them. In Section 4, we consider generalized Killing and twistor spinors and find their description in terms of differential forms of the exterior bundle on the ambient manifold. Section 5 includes the description of generalized Killing vectors in terms of generalized connection and definition of generalized KY forms. In Section 6, relation between generalized Killing spinors and generalized KY forms is investigated. Section 7 concludes the paper.
@@ II Generalized Geometry
or we can also denote it in the matrix form as $\mathcal{X}=\left(\begin{array}[]{cc}X\\ \xi\end{array}\right)$. One can define a bilinear form <,><\,,\,> on $E$ which is written for generalized vectors $\mathcal{X}=X+\xi$ and $\mathcal{Y}=Y+\eta$ as follows
$\mathcal{X},\mathcal{Y}>=\frac{1}{2}\big(i_{X}\eta+i_{Y}\xi\big)$ (1)
where $i_{X}$ denotes the interior derivative or contraction with respect to the vector field $X$. So, for a single generalized vector, we have $\mathcal{X},\mathcal{X}>=i_{X}\xi$. This bilinear form has signature $(n,n)$ and with the orientation defined for $E$, the structure group of $E$ reduces to $SO(n,n)$ [2]. The invariance of <,><\,,\,> under $SO(n,n)$ gives shear transformations for generalized vectors. For a 2-form $B\in\Lambda^{2}M$ and a 2-vector $\beta\in\Lambda_{2}M$, we have $B$- and $\beta$-transforms of $\mathcal{X}$ as follows [2]
@@
which is the generalized metric induced by $g$ and $B$. Since $\mathcal{G}^{2}=I$, $\pm 1$ eigenspaces of $\mathcal{G}$ which are denoted by $V_{\pm}$ give a metric splitting of $E$
@@
From the basis vectors $\{X_{a}\}\in\Gamma TM$ and basis 1-forms $\{e^{a}\}\in\Gamma T^{*}M$ for a=1,,na=1,...,n, we can define a basis for generalized vectors as
@@
where $\epsilon=\pm 1$ for $V_{\pm}$ and A=1,,2nA=1,...,2n. If the generalized vector basis $\mathcal{X}_{A}\in V_{+}$ then $\epsilon=+1$ and if $\mathcal{X}_{A}\in V_{-}$ then $\epsilon=-1$. Since $A$ takes values of 1,…,2n and $a$ takes values of 1,…,n, the above equation is equivalent to
$\mathcal{X}_{A}=\{X_{a}+e^{a},X_{n+a}-e^{n+a}\}$
and the left hand side and right hand side indices are equivalent in number and the index $A=\{a,n+a\}$ is just a nomenclature of the right hand side. $\mathcal{G}$-duals of $\mathcal{X}_{A}$ correspond to basis for generalized 1-forms
@@
Similar to the Lie bracket [,][\,,\,] of vector fields on $\Gamma TM$, we can also define a bracket operation on $E$ called Courant bracket as $_{C}:\Gamma E\times\Gamma E\rightarrow\Gamma E$. For two generalized vectors $\mathcal{X}=X+\xi$ and $\mathcal{Y}=Y+\eta$, Courant bracket is defined as
$\mathcal{X},\mathcal{Y}]_{C}=[X,Y]+\mathcal{L}_{X}\eta-\mathcal{L}_{Y}\xi-\frac{1}{2}d\left(i_{X}\eta-i_{Y}\xi\right)$ (7)
@@
for $X,Y,Z\in\Gamma TM$. The cohomology class of $H$ classifies the Courant algebroids on $M$ up to isomorphism [22, 23, 24]. In the presence of non-zero $H$, one can modify the Courant bracket and define the twisted Courant bracket $_{H}$ as follows
$\mathcal{X},\mathcal{Y}]_{H}=[\mathcal{X},\mathcal{Y}]_{C}-i_{X}i_{Y}H.$ (10)
$_{H}$ also defines a Courant algebroid structure for $dH=0$. Another bracket operation on $E$ which satisfies the Jacobi identity but is not antisymmetric is the Dorfman bracket $_{D}$ and it is defined as
$\mathcal{X},\mathcal{Y}]_{D}=[X,Y]+\mathcal{L}_{X}\eta-i_{Y}d\xi.$ (11)
Courant bracket corresponds to the antisymmetrization of the Dorfman bracket and it can also be written as
$\mathcal{X},\mathcal{Y}]_{C}=[\mathcal{X},\mathcal{Y}]_{D}-d<\mathcal{X},\mathcal{Y}>.$ (12)
@@
$\mathbb{L}_{\mathcal{X}},\mathbb{L}_{\mathcal{Y}}]=\mathbb{L}_{[\mathcal{X},\mathcal{Y}]_{D}}$ (13)
@@
We can define a generalized connection $\mathbb{D}$ on $E$ which is compatible with the bilinear form <,><\,,\,> corresponding to the linear operator
@@
However, there is a freedom in choosing a connection since $\mathbb{D}^{\prime}=\mathbb{D}+\alpha$ also defines a connection where $\alpha\in\Gamma(E^{*}\otimes o(E))$ and $o(E)$ denotes the bundle of skew-symmetric endomorphisms of $E$ w.r.t. <,><\,,\,> [26, 27]. The generalized torsion 3-form $T_{\mathbb{D}}\in\Gamma(\Lambda^{3}E^{*})$ of a generalized connection $\mathbb{D}$ on $E$ is defined by
@@
In the presence of a non-zero 3-form field $H$, the conditions of metric compatibility and vanishing generalized torsion do not uniquely fix the connection; rather, they define a class of admissible connections. However, as discussed in [15] (see Eq. 3.4 therein), the relevant physical operators acting on spinors are defined intrinsically by the generalized exterior derivative. Therefore, the resulting equations are independent of the specific choice of connection within this class. Consequently, for explicit calculations, we are free to adopt the canonical choice of the generalized connection [26] without loss of generality:
@@ III Inner product classes and Properties of Generalized Spinor Bilinears
which is isomorphic to $\mathbb{R}^{2^{n-1}}\oplus\mathbb{R}^{2^{n-1}}$. The spinor inner product $(\,,\,):S\times S\rightarrow\mathbb{R}$ of two generalized geometry spinors $\phi$ and $\psi$ is written as
@@
where $J$ is any matrix that we choose as identity matrix, $\mathcal{J}$ is an inner automorphism of the Clifford algebra and $(\,\,)_{n}$ denotes the projection to the $n$-form part. For $\mathcal{J}$, we have two choices which are $\xi$ and $\xi\eta$ automorphisms where $\eta$ acts on a $p$-form $\omega$ as $\omega^{\eta}=(-1)^{p}\omega$ and $\xi$ acts as $\omega^{\xi}=(-1)^{\lfloor\frac{p}{2}\rfloor}\omega$ with $\xi\eta$ is the composition of both automorphisms. Here, $\lfloor\rfloor$ denotes the floor function. We can determine the symmetry properties and automorphism classes of inner products for different values of $n$ by using the spinor inner product classification tables given in [28, 29] as follows
@@
n(mod 4)n($\text${mod }4) 0 -$\mathbb{R}\text{-swap }\xi$ and -$\mathbb{R}\text{-sym}\oplus\mathbb{R}\text{-sym }\xi\eta$ 11 -$\mathbb{R}\text{-sym}\oplus\mathbb{R}\text{-sym }\xi$ and -$\mathbb{R}\text{-swap }\xi\eta$ 22 -$\mathbb{R}\text{-swap }\xi$ and -$\mathbb{R}\text{-skew}\oplus\mathbb{R}\text{-skew }\xi\eta$ 33 -$\mathbb{R}\text{-skew}\oplus\mathbb{R}\text{-skew }\xi$ and -$\mathbb{R}\text{-swap }\xi\eta$
Since the spinor space is real, we have real inner products, sym and skew denotes the symmetric and skew inner products and swap means that inner product changes the chiral elements to each other, $\xi$ and $\xi\eta$ correspond to the relevant automorphism class. So, we have for n=0,1(mod 4)n=0,1($\text${mod }4) symmetric and for n=2,3(mod 4)n=2,3($\text${mod }4) skew inner products. Since, only the $n$-form part remains in the spinor inner product, one can see that, for n=0(mod 2)n=0($\text${mod }2), inner products of elements
S+×S and S×S+ are zero $^{+}\times S^{-}\text${ and $^{-}\times S^{+}\text${ are zero }
S+×S+ and S×S are non-zero $^{+}\times S^{+}\text${ and $^{-}\times S^{-}\text${ are non-zero }
since the inner products of opposite chirality spinors cannot include even degree part and for n=1(mod 2)n=1($\text${mod }2), inner products of elements
S+×S+ and S×S are zero $^{+}\times S^{+}\text${ and $^{-}\times S^{-}\text${ are zero }
S+×S and S×S+ are non-zero $^{+}\times S^{-}\text${ and $^{-}\times S^{+}\text${ are non-zero }
since the inner products of same chirality spinors cannot include odd degree part. Then, for n=0(mod 2)n=0($\text${mod }2), we need to choose -$\mathbb{R}\text{-sym}\oplus\mathbb{R}\text{-sym}$ or -$\mathbb{R}\text{-skew}\oplus\mathbb{R}\text{-skew}$ inner products since only the inner products of same chirality spinors are non-zero in this case and for n=1(mod 2)n=1($\text${mod }2), we need to choose -$\mathbb{R}\text{-swap}$ inner products since only the inner products of opposite chirality spinors are non-zero in this case. As a result, we have the inner product symmetries and automorphisms classes as follows
@@
n(mod 4)n($\text${mod }4) 0 -$\mathbb{R}\text{-sym}\oplus\mathbb{R}\text{-sym }\xi\eta$ 11 -$\mathbb{R}\text{-swap (sym) }\xi\eta$ 22 -$\mathbb{R}\text{-skew}\oplus\mathbb{R}\text{-skew }\xi\eta$ 33 -$\mathbb{R}\text{-swap (skew) }\xi\eta$
Hence, we can write the spinor inner product for generalized spinors as
@@
which corresponds to the Mukai pairing of forms. It can be seen from the previous analysis of inner product classes that the automorphism class must be taken as $\xi\eta$ to have a non-zero spinor inner product. For any two generalized spinors $\phi$, $\psi$ and a generalized form $\mathcal{\alpha}$, we can write
@@
where the action of a generalized form on a generalized geometry spinor $\alpha.\psi$ can be defined inductively from the action of a generalized vector on a differential form given in (14) . From the symmetry properties of the inner product and the choice of the automorphism class as $\xi\eta$, we have
@@
For a spinor $\phi$ and a $p$-form $\alpha_{p}$, we can define the spinor bilinear $p$-form as $(\phi,\alpha_{p}.\phi)$ and writing the spinor bilinears for a generalized geometry spinor $\phi$ and a generalized $p$-form $\mathcal{\alpha}_{p}$, we have
@@
where we have used $\mathcal{\alpha}_{p}^{\xi\eta}=(-1)^{p(p-1)/2}(-1)^{p}\mathcal{\alpha}_{p}$. Moreover, for $n$ even, $p$ must be even, and for $n$ odd, $p$ must be odd, since for $n$ even $S^{\pm}\times S^{\pm}$ and for $n$ odd $S^{\pm}\times S^{\mp}$ inner products are non-zero. For a definite chirality spinor $\phi$, since the action of a generalized vector on a generalized geometry spinor changes the chirality of the spinor, $\alpha_{p}.\phi$ has the same chirality if $p$ is even and has the reverse chirality if $p$ is odd. Then, by considering those properties, $(-1)^{\frac{n(n-1)+p(p+1)}{2}}=1$ only for n+p=0(mod 4)n+p=0($\text${mod }4) and
@@
is non-zero and for other cases the bilinear form is zero. In general, for two different generalized geometry spinors $\phi$ and $\psi$ $(\phi\neq\psi)$, we have
@@
and other cases are zero. Then, only some special degree spinor bilinears are non-zero depending on $n$ and $p$. For two generalized geometry spinors $\phi$ and $\psi$, we can write the spinor bilinear as the tensor product of a dual spinor and a spinor which can be written as a sum of $p$-form bilinears in terms of generalized $p$-form basis as